# Unlock the Power of Algorithm Features: A Generalization Analysis for Algorithm Selection

### v1

@@ Abstract
In the field of algorithm selection research, the discussion surrounding algorithm features has been significantly overshadowed by the emphasis on problem features. Although a few empirical studies have yielded evidence regarding the effectiveness of algorithm features, the potential benefits of incorporating algorithm features into algorithm selection models and their suitability for different scenarios remain unclear. It is evident that relying solely on empirical research cannot adequately elucidate the mechanisms underlying performance variations. In this paper, we address this gap by proposing the first provable guarantee for algorithm selection based on algorithm features, taking a generalization perspective. We analyze the benefits and costs associated with algorithm features and investigate how the generalization error is affected by several factors. Specifically, we examine adaptive and predefined algorithm features under transductive and inductive learning paradigms, respectively, and derive upper bounds for the generalization error based on their model’s Rademacher complexity. Our theoretical findings not only provide tight upper bounds, but also offer analytical insights into the impact of various factors, including model complexity, the number of problem instances and candidate algorithms, model parameters and feature values, and distributional differences between the training and test sets. Notably, we demonstrate that algorithm feature-based models outperform traditional models relying solely on problem features in complex multi-algorithm scenarios in terms of generalization, and are particularly well-suited for deployment in scenarios under distribution shifts, where the generalization error exhibits a positive correlation with the chi-square distance between training and test sets.
@@ 1 Introduction
For most computational problems, it is widely recognized that no single algorithm outperforms all others on every problem instance [1]. This phenomenon, known as performance complementarity, has been observed across various domains, including machine learning tasks [2, 3], optimization problems [4, 5], NP-hard problems [6, 7], and so on. To address this challenge, per-instance algorithm selection has been extensively studied, which aims to determine the algorithm that is expected to perform best on a specific problem instance [8]. The existing algorithm selection techniques can be categorized into two groups based on their consideration of algorithm features: approaches based on problem features and approaches based on both problem and algorithm features (referred to as problem feature-based methods and algorithm feature-based methods hereafter for brevity). Fig. 1 illustrates the distinctions between these two methodologies. Due to disparities in information sources, problem feature-based methods directly maps the problem feature space to the algorithm space, while the algorithm feature-based methods explore matching patterns between problem features and algorithm features to determine the most suitable algorithm.
Many problem feature-based methods are primarily grounded in the framework proposed by Rice [9], which conceptualizes algorithm selection as a mapping from problem features to algorithms, without taking into account algorithm features. These methods often formulate the algorithm selection task as either a performance regression or a multi-class classification problem. Specifically, regression-based techniques [10] directly predict the performance of each algorithm based on the problem feature set. Conversely, multi-class classification approaches [11, 12] assign scores to algorithms based on their relative suitability for a specific problem and select the most appropriate algorithm based on these scores. In addition, several other methodologies have been proposed. For instance, some studies formalize algorithm selection as a collaborative filtering problem [13, 14] and utilize a sparse matrix with performance data available for only a few algorithms on each problem instance. And similarity-based methods [15, 16] select algorithms based on the similarity between the current problem instance and previously encountered instances. There are also some hybrid methods [17, 18] combine multiple techniques above.
@@
While these studies have provided some evidence of the positive impact of algorithm features on the algorithm selection task, they are still insufficient to offer compelling guidance for the application of algorithm features in real-world solutions. On one hand, within the limited scope of empirical research on algorithm features, the comparison between models incorporating algorithm features and models utilizing only problem features has yielded mixed results in different scenarios, as evidenced by studies such as [22, 24]. The conditions under which algorithm features should be employed, as well as the specific types of algorithm features to consider, remain unclear. On the other hand, several factors, including the scale of training data and the extent of distributional shifts, can significantly influence the design of algorithm selection approaches. The rigorous analysis of how to incorporate these factors into the decision-making process of algorithm feature utilization is currently lacking. Consequently, the use of algorithm features remains primarily at the stage of experimental validation, with ongoing debates regarding their effectiveness and appropriate application conditions [26] for practical deployment.
@@
Intuitively, considering algorithm features in the context of algorithm selection may enhance the predictive capacity of models by introducing additional information. If these features prove informative and there is sufficient training data, the model hold the potential for achieving better performance on testing data, as demonstrated in the results presented in [24]. However, larger-capacity models necessitate a larger quantity of training data, which may pose challenges in scenarios where collecting performance data for candidate algorithms is prohibitively expensive. Notably, algorithm features appear to improve the model’s ability to generalize to changed distribution, encompassing not only diverse problem instances but also novel algorithms that were not encountered during training. The degree to which the model benefits from algorithm features in terms of inductive generalization is contingent upon various factors, including the model’s theoretical complexity, the scale of the problem instances and candidate algorithms involved, the property of networks and loss functions, as well as the distribution variation, among others. This paper aims to explore the generalization nature of algorithm feature-based models and provide a theoretical analysis of how the interplay of these factors influences the model’s generalization performance. By this means, we aim to elucidate the advantage and trade-off associated with algorithm features.
@@
Before the incorporation of algorithm features, existing methods are predominantly designed within the transductive learning paradigm [27], with little deliberate consideration given to the theoretical generalization error of these approaches. While not explicitly stated in these studies, most of them assumed the training and test sets shared the same distribution [^1] This statement is limited to the algorithm selection on static scenarios only. It is worth noting that some algorithm selection methods for data stream [28, 29] may take into account distribution shifts in the problem space [30]. However, the scope of discussion in this paper is limited to static data. and trained models solely on problem features. Once algorithm features are considered, both adaptive features and predefined features will impart the model with diverse aspects of generalization capabilities across different learning paradigms. Adaptive features are derived based on the adaptability between algorithms and problems within the training data, making them unable to generalize to new candidate algorithms. Conversely, predefined features directly capture the intrinsic properties of algorithms, enabling them to exhibit generalization performance even in scenarios with distribution shifts. Therefore, we will explore the model generalization under two paradigms, i.e., transductive [27] and inductive [31] learning settings. To ensure the broad applicability of our theoretical research, we adopt a higher level of abstraction for modeling, i.e., the multilayer perceptrons with arbitrary depths and hidden layer designs, employing 1-Lipschitz activation functions. Under both learning paradigms, we provide rigorous upper bounds of the Rademacher complexity and generalization error for models based on adaptive algorithm features, predefined algorithm features, as well as regression and multi-class classification models based on problem features in Sections 3 and 4. Based on these theoretical investigations, we draw the following conclusions to inspire the practical application of algorithm features in real-world scenarios:
@@ 2.1 Model Definition and Basic Notations
To discuss the impact of algorithm features, we first define the algorithm selection model based on different types of algorithm features. In this paper, we discuss two types of algorithm features, corresponding to different learning mechanism, namely adaptive features and predefined features. As shown in Fig. 2, predefined features, derived from a automatic or manual extraction process, can capture inherent algorithmic properties, e.g., features from algorithmic code or algorithm description. On the other hand, adaptive features learned through matching relationships between algorithms and problem instances (achieved by embedding layer), provide nuanced insights into algorithmic suitability for specific problem domains.
Figure 2: Illustration of algorithm selection model based on the adaptive algorithm features and the predefined algorithm features.
@@
We provide the notations in the following. Given a dataset $\mathcal{D}$, consisting of a problem set $\mathcal{P}$, a algorithm set $\mathcal{A}$ for solving problem instances in $\mathcal{P}$, and a performance metric $\mathcal{M}$: $\mathcal{P}\times\mathcal{A}\rightarrow\mathbb{R}$ which quantifies the performance of any algorithm $A_{j}\in\mathcal{A}$ on each problem instance $P_{i}\in\mathcal{P}$, algorithm feature-based algorithm selection model should construct a selector $\mathbf{f}:\mathcal{P}\times\mathcal{A}\rightarrow\mathbb{R}$ or $\mathbf{f}:\mathcal{P}\times\mathcal{A}\rightarrow\{0,1\}$ that determines the compatibility (or degree of compatibility) between any algorithm and any problem instance. The dataset $\mathcal{D}$ is divided into a training set and a test set, represented as $\mathcal{S}$ and $\mathcal{T}$, respectively. Their sizes are denoted as $|\mathcal{S}|$ and $|\mathcal{T}|$. In methods that do not consider algorithm features, most studies require learning the mapping from $\mathcal{P}$ to $\mathcal{A}$, such as algorithm selection approaches based on multi-class classification methods of performance regression methods. In this case, the size of the training data is equal to the number of problem instances. When the role of algorithm features is fully considered, it is necessary to define the sets of both problems and algorithms for the training and test purpose, denoted as $\mathcal{S}=\{\mathcal{S}_{\mathcal{P}}\cup\mathcal{S}_{\mathcal{A}}\cup\mathcal{S}_{\mathcal{M}}\}$ and $\mathcal{T}=\{\mathcal{T}_{\mathcal{P}}\cup\mathcal{T}_{\mathcal{A}}\cup\mathcal{T}_{\mathcal{M}}\}$, respectively. For the rest of this paper, unless otherwise specified, the candidate algorithms in $\mathcal{S}_{\mathcal{A}}$ and $\mathcal{T}_{\mathcal{A}}$ are assumed to be the same, i.e., $\mathcal{S}_{\mathcal{A}}=\mathcal{T}_{\mathcal{A}}$. Moreover, since the model’s input is a problem-algorithm pair, the size of the training data and test data corresponds to $|\mathcal{S}|=|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|$ and $|\mathcal{T}|=|\mathcal{T}_{\mathcal{P}}|\cdot|\mathcal{T}_{\mathcal{A}}|$, respectively.
@@
In the following theoretical analysis, we will discuss a simple yet sufficiently universal and effective model. The input will be fed into a deep neural network $\mathbf{f}$ with $l$ layers, and the parameters in each layer are denoted as $W^{(i)}(i\in\{1,2,\cdots,l\})$. Any 1-Lipschitz activation functions (e.g., ReLU, Leaky ReLU, SoftPlus, Tanh, Sigmoid, ArcTan, or Softsign) can be used. For model in Fig. 2(a), the input includes the problem features and the algorithm index. For ease of modeling, the algorithm features in the embedding layer are combined with problem features, which are collectively represented as the parameters of the first layer of the deep network, denoted as $W^{(0)}=\begin{bmatrix}PF_{1},PF_{1},\cdots\\ AF_{1},AF_{2},\cdots\end{bmatrix}$, where $PF_{i}$ and $AF_{j}$ are frozen problem features and optimized algorithm features. Correspondingly, the input is a one-hot column vector, where non-zero positions signify the indices of the problem and algorithm. The main body of the model remains consistent with the previous definition. We denote the model based on adaptive features as Modela. For model in Fig. 2(b), the input includes the problem features and the predefined algorithm features. The model body is the same as the previously defined model. We denote the model based on predefined features as Modelb.
@@ 2.2 Transductive and Inductive Generalization in Algorithm Selection
^$\widehat{\mathfrak{R}}_{n}(\mathcal{V})\triangleq\frac{1}{n}\mathbf{E}_{\bm{\sigma}}\left\{\sup_{v\in\mathcal{V}}\sum_{i=1}^{n}\sigma_{i}v\left(x_{i}\right)\right\}$ (3)
@@ 3 Generalization Error of the Adaptive Feature-based Model
Compared to using only problem features, incorporating both problem and algorithm features undoubtedly increases the model’s capacity, capturing more comprehensive information. In Model a, an embedding layer is introduced as the first layer of the model to represent adaptive features, where each column vector corresponds to the representation of a candidate algorithm. This section will analyze the theoretical performance based on Modela.
According to Definition 1, the transductive Rademacher Complexity can be calculated as:
@@
$\left(\frac{1}{|\mathcal{S}|}+\frac{1}{|\mathcal{T}|}\right)\mathbb{E}_{\bm{\sigma}}\left[\sup\sum_{x_{i}\in\mathcal{S}}\gamma\sigma_{i}\mathbf{f}(x_{i})\right]=\gamma\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{F},p)$
@@
Follow the same definition in Theorem 1. Let $x_{i}\in\mathcal{S}$ denote any training instance with problem feature vector $\textbf{\emph{PF}}_{i}$ and algorithm feature vector $\textbf{\emph{AF}}_{i}$ in embedding layer. Then the transductive Rademacher complexity of Modela, $\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{L},p)$, is bounded by:
$\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{L},p)\leq\frac{\gamma L(|\mathcal{S}|+|\mathcal{T}|)}{|\mathcal{S}||\mathcal{T}|}\sup_{x_{i}\in\mathcal{S}}(\|[\textbf{\emph{PF}}_{i}\|_{2}+\|[\textbf{\emph{AF}}_{i}\|_{2})$ (17)
@@
Let $\mathcal{D}\rightarrow\mathcal{S}\cup\mathcal{T}$ is a partition of random permutation of $\mathcal{D}$ with partition ratio $\eta=\frac{|\mathcal{T}|}{|\mathcal{S}|}<1$  [^2] The training set scale is larger than the test set, as a research convention., and $\mathcal{S}_{\mathcal{A}}$, $\mathcal{S}_{\mathcal{P}}$, $\mathcal{T}_{\mathcal{A}}$, $\mathcal{T}_{\mathcal{P}}$, denote the candidate algorithms and problems in $\mathcal{S}$ and $\mathcal{T}$. $L$ is the Lipschitz constants of the model excluding algorithm embedding layer, $\gamma$ is the Lipschitz constant of the loss function, and $W^{(0)}$ is the parameters composed of the problems features and parameters in algorithm embedding layer. $Error_{\mathcal{T}}(\mathbf{f})$ and $Error_{\mathcal{S}}(\mathbf{f})$ denote the test error and empirical error of Modela, and $c_{0}\triangleq\sqrt{\frac{32\ln(4e)}{3}}$. Then with probability of at least $1-\delta$, for all $\ell\left(\mathbf{f}(x_{i}),y_{i}\right)\in\mathcal{L}$,
@@
According to Theorem 2, we observe that the generalization error under the usage of adaptive features is related to three factors: the training data $|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|$, the partition ratio $\eta$, and some parameters in model including $L$, $W^{(0)}$, and $\gamma$. When the size of the training set is large enough such that $|\mathcal{S}|\gg\frac{\gamma^{2}L^{2}\|W^{(0)}\|^{2}_{2}(1+\eta)^{2}}{c^{2}\eta}$ where $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$, the trend of the generalization error tends to decrease continuously at a rate of $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$ as the training scale grows. While if the training data is insufficient, the model parameters and the size of the training set will jointly affect the generalization error, following the slack term $\mathcal{O}\left(\frac{\gamma L\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\right)$. This is because the model may fail to capture all the patterns in the data, leading to underfitting. In such cases, it is not advisable to have too many or overly complex model parameters. Based on Theorem 2, it can also be observed that the generalization error is relevant to the product of the norms of each layer’s parameter matrix, as determined by the computation of the upper bound for the Lipschitz constant $L$. Because the $L_{2}$-norm emphasizes the maximum singular value of the parameter matrix [36], outliers in parameter matrix may have a significant impact on the maximum singular value. This highlights the importance of employing preprocessing techniques such as normalization and outlier detection in practical usage to mitigate the occurrence of outliers in the parameter matrices of neural networks. In our subsequent research within the inductive setting, we will derive a Frobenius norm with improved properties for bounding generalization error, which will demonstrate enhanced resilience to outliers [37], providing better control over the upper bound of generalization error.
@@
On the other hand, it is not difficult to observe from Theorem 2 that, when using the adaptive algorithm feature, the increase in the number of candidate algorithms will have a positive impact on reducing generalization error, following the slack term $\mathcal{O}\left(\sqrt{\frac{1}{|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$, which is significantly different from traditional learning approaches that focus on learning the mapping from problems to algorithms. When the algorithm feature is absence, for example, some studies model algorithm selection as a multi-class classification problem or performance regression problem based on problem features. In this case, the derivation process is similar to Theorem 2, which allows us to easily obtain upper bounds for the transductive generalization error of the model under the performance regression approach (MLPr) and the multi-class classification approach (MLPc). We present them in the form of Corollary 2:

### v2

@@ Abstract
In the algorithm selection research, the discussion surrounding algorithm features has been significantly overshadowed by the emphasis on problem features. Although a few empirical studies have yielded evidence regarding the effectiveness of algorithm features, the potential benefits of incorporating algorithm features into algorithm selection models and their suitability for different scenarios remain unclear. In this paper, we address this gap by proposing the first provable guarantee for algorithm selection based on algorithm features, taking a generalization perspective. We analyze the benefits and costs associated with algorithm features and investigate how the generalization error is affected by different factors. Specifically, we examine adaptive and predefined algorithm features under transductive and inductive learning paradigms, respectively, and derive upper bounds for the generalization error based on their model’s Rademacher complexity. Our theoretical findings not only provide tight upper bounds, but also offer analytical insights into the impact of various factors, such as the training scale of problem instances and candidate algorithms, model parameters, feature values, and distributional differences between the training and test data. Notably, we demonstrate how models will benefit from algorithm features in complex scenarios involving many algorithms, and proves the positive correlation between generalization error bound and $\chi^{2}$-divergence of distributions.
@@ 1 Introduction
For most computational problems, it is widely recognized that no single algorithm outperforms all others on every problem instance [1]. This phenomenon, known as performance complementarity, has been observed across various domains, including machine learning tasks [2, 3], optimization problems [4, 5], NP-hard problems [6, 7], and so on. To address this challenge, per-instance algorithm selection has been extensively studied, which aims to determine the algorithm that is expected to perform best on a specific problem instance [8]. Existing algorithm selection techniques can be categorized into two groups based on their consideration of algorithm features: approaches based on problem features and approaches based on both problem and algorithm features (referred to as problem feature-based methods and algorithm feature-based methods hereafter for brevity). Fig. 1 illustrates the distinctions between these two methodologies.
Many problem feature-based methods are primarily grounded in the framework proposed by Rice [9], which conceptualizes algorithm selection as a mapping from problem features to algorithms, without taking into account algorithm features. These methods often formulate the algorithm selection task as either a performance regression or a multi-class classification problem. Specifically, regression-based techniques [10] directly predict the performance of each algorithm based on the problem feature set. Conversely, multi-class classification approaches [11, 12, 13] assign scores to algorithms based on their relative suitability for a specific problem and select the most appropriate algorithm based on these scores. In addition, several other methodologies have been proposed. For instance, some studies formalize algorithm selection as a collaborative filtering problem [14, 15] and utilize a sparse matrix with performance data available for only a few algorithms on each problem instance. And similarity-based methods [16, 17] select algorithms based on the similarity between the current problem instance and previously encountered instances. There are also some hybrid methods [18, 19] combine multiple techniques above.
@@
While these studies have provided some evidence of the positive impact of algorithm features on the algorithm selection task, they are still insufficient to offer compelling guidance for the application of algorithm features in real-world solutions. On one hand, within the limited scope of empirical research on algorithm features, the comparison between models incorporating algorithm features and models utilizing only problem features has yielded mixed results in different scenarios, as evidenced by studies such as [23, 25]. The conditions under which algorithm features should be employed, as well as the specific types of algorithm features to consider, remain unclear. On the other hand, several factors, including the scale of training data and the extent of distributional shifts, can significantly influence the design of algorithm selection approaches. The rigorous analysis of how to incorporate these factors into the decision-making process of algorithm feature utilization is currently lacking. Consequently, the use of algorithm features remains primarily at the stage of experimental validation, with ongoing debates regarding their effectiveness and appropriate application conditions for practical deployment.
@@
Intuitively, considering algorithm features in the context of algorithm selection may enhance the predictive capacity of models by introducing additional information. If these features prove informative and there is sufficient training data, the model hold the potential for achieving better performance on testing data [25]. However, larger-capacity models necessitate a larger quantity of training data, which may pose challenges in scenarios where collecting performance data for candidate algorithms is prohibitively expensive. Notably, algorithm features appear to improve the model’s ability to generalize to changed distribution, encompassing not only diverse problem instances but also novel algorithms that were not encountered during training. The degree to which the model benefits from algorithm features in terms of inductive generalization is contingent upon various factors, including the model’s theoretical complexity, the scale of the problem instances and candidate algorithms involved, the property of networks and loss functions, as well as the distribution variation, among others. This paper aims to explore the generalization nature of algorithm feature-based models and provide a theoretical analysis of how the interplay of these factors influences the model’s generalization performance. By this means, we aim to elucidate the advantage and trade-off associated with algorithm features.
@@
Before the incorporation of algorithm features, existing methods are predominantly designed within the transductive learning paradigm [27], with little deliberate consideration given to the theoretical generalization error of these approaches. While not explicitly stated in these studies, most of them assumed the training and test sets shared the same distribution [^1] This statement is limited to the algorithm selection on static scenarios only. It is worth noting that some existing algorithm selection methods for streaming data [28, 29] may take into account distribution changes in the problem stream [30]. However, the scope of discussion in this paper is limited to static data. and trained models solely on problem features. Once algorithm features are considered, both adaptive features and predefined features will impart the model with diverse aspects of generalization capabilities across different learning paradigms. Adaptive features are derived based on the adaptability between algorithms and problems within the training data, making them unable to generalize to new candidate algorithms. Conversely, predefined features directly capture the intrinsic properties of algorithms, enabling them to exhibit generalization performance even in scenarios with distribution shifts. Therefore, we will explore the model generalization under two paradigms, i.e., transductive [27] and inductive [31] learning settings. To ensure the broad applicability of our theoretical research, we adopt a higher level of abstraction for modeling, i.e., the multilayer neural network with arbitrary depths and hidden layer designs, employing 1-Lipschitz activation functions. Under both learning paradigms, we provide the upper bounds of the Rademacher complexity and generalization error for models based on adaptive algorithm features, predefined algorithm features, as well as regression and multi-class classification models based on problem features in Sections 3 and 4. Based on these theoretical investigations, we draw the following conclusions to inspire the practical application of algorithm features in real-world scenarios:
@@ 2.1 Model Definition and Basic Notations
To discuss the impact of algorithm features, we first define the algorithm selection model based on different types of algorithm features. In this paper, we discuss two types of algorithm features, corresponding to different learning mechanism, namely adaptive features and predefined features. Predefined features, derived from a automatic or manual extraction process, can capture inherent algorithmic properties, e.g., features from algorithmic code or algorithm description. On the other hand, adaptive features learned through matching relationships between algorithms and problem instances (achieved by embedding layer), provide nuanced insights into algorithmic suitability for specific problem domains.
@@
We provide the notations in the following. Given a dataset $\mathcal{D}$, consisting of a problem set $\mathcal{P}$, a algorithm set $\mathcal{A}$ for solving problem instances in $\mathcal{P}$, and a performance metric $\mathcal{M}$: $\mathcal{P}\times\mathcal{A}\rightarrow\mathbb{R}$ which quantifies the performance of any algorithm $A_{j}\in\mathcal{A}$ on each problem instance $P_{i}\in\mathcal{P}$, algorithm feature-based algorithm selection model should construct a selector $\mathbf{f}:\mathcal{P}\times\mathcal{A}\rightarrow\mathbb{R}$ or $\mathbf{f}:\mathcal{P}\times\mathcal{A}\rightarrow\{0,1\}$ that determines the compatibility (or degree of compatibility) between any algorithm and any problem instance. The dataset $\mathcal{D}$ is divided into a training set and a test set, represented as $\mathcal{S}$ and $\mathcal{T}$, respectively. Their sizes are denoted as $|\mathcal{S}|$ and $|\mathcal{T}|$. In methods that do not consider algorithm features, most studies require learning the mapping from $\mathcal{P}$ to $\mathcal{A}$, such as algorithm selection approaches based on multi-class classification methods or performance regression methods. In this case, the size of the training data is equal to the number of problem instances. When the role of algorithm features is fully considered, it is necessary to define the sets of both problems and algorithms for the training and test purpose, denoted as $\mathcal{S}=\{\mathcal{S}_{\mathcal{P}}\cup\mathcal{S}_{\mathcal{A}}\}$ and $\mathcal{T}=\{\mathcal{T}_{\mathcal{P}}\cup\mathcal{T}_{\mathcal{A}}\}$, respectively. For the rest of this paper, unless otherwise specified, the candidate algorithms in $\mathcal{S}_{\mathcal{A}}$ and $\mathcal{T}_{\mathcal{A}}$ are assumed to be the same, i.e., $\mathcal{S}_{\mathcal{A}}=\mathcal{T}_{\mathcal{A}}$. Moreover, since the model’s input is a problem-algorithm pair, the size of the training data and test data corresponds to $|\mathcal{S}|=|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|$ and $|\mathcal{T}|=|\mathcal{T}_{\mathcal{P}}|\cdot|\mathcal{T}_{\mathcal{A}}|$, respectively.
@@
In the following theoretical analysis, we will discuss a simple yet sufficiently universal and effective model. The input will be fed into a deep neural network $\mathbf{f}$ with $l$ layers, and the parameters in each layer are denoted as $W^{(i)}(i\in\{1,2,\cdots,l\})$. Any 1-Lipschitz activation functions (e.g., ReLU, Leaky ReLU, SoftPlus, Tanh, Sigmoid, ArcTan, or Softsign) can be used. For model based on adaptive features, the input includes the problem features and the algorithm index. For ease of modeling, the algorithm features in the embedding layer are combined with problem features, which are collectively represented as the parameters of the first layer of the deep network, denoted as $W^{(0)}=\begin{bmatrix}PF_{1},PF_{1},\cdots\\ AF_{1},AF_{2},\cdots\end{bmatrix}$, where $PF_{i}$ and $AF_{j}$ are frozen problem features and optimized algorithm features. Correspondingly, the input is a one-hot column vector, where non-zero positions signify the indices of the problem and algorithm. The main body of the model remains consistent with the previous definition. We denote the model based on adaptive features as Modela. For model based on predefined features, the input includes the problem features and the predefined algorithm features. The model body is the same as the previously defined model. We denote the model based on predefined features as Modelb.
@@ 2.2 Transductive and Inductive Generalization in Algorithm Selection
where $\mathbf{E}$ denotes the mathematical expectation in this paper.
@@
^$\widehat{\mathfrak{R}}_{n}(\mathcal{V})\triangleq\frac{1}{n}\mathbf{E}_{\bm{\sigma}}\left\{\sup_{\mathbf{v}\in\mathcal{V}}\sum_{i=1}^{n}\sigma_{i}v\left(x_{i}\right)\right\}$ (3)
@@
In the following, we will analyze the upper bound of the generalization error based on these definitions.
@@ 3 Generalization Error of the Adaptive Feature-based Model
Compared to using only problem features, incorporating both problem and algorithm features usually increases the model’s capacity, capturing more comprehensive information. This section analyzes the theoretical performance based on Modela. According to Definition 1, the transductive Rademacher Complexity can be calculated as:
@@
$\left(\frac{1}{|\mathcal{S}|}+\frac{1}{|\mathcal{T}|}\right)\mathbf{E}_{\bm{\sigma}}\left[\sup\sum_{x_{i}\in\mathcal{S}}\gamma\sigma_{i}\mathbf{f}(x_{i})\right]=\gamma\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{F},p)$
@@
Follow the same definition in Theorem 1. Let $x_{i}\in\mathcal{S}$ denote any training instance with problem feature vector $PF_{i}$ and algorithm feature vector $AF_{i}$ in embedding layer. Then the transductive Rademacher complexity of Modela, $\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{L},p)$, is bounded by:
$\mathfrak{R}_{|\mathcal{S}|+|\mathcal{T}|}(\mathcal{L},p)\leq\frac{\gamma L(|\mathcal{S}|+|\mathcal{T}|)}{|\mathcal{S}||\mathcal{T}|}\sup_{x_{i}\in\mathcal{S}}(\|[PF_{i}\|_{2}+\|[AF_{i}\|_{2})$ (17)
@@
Let $\mathcal{D}\rightarrow\mathcal{S}\cup\mathcal{T}$ is a partition of random permutation of $\mathcal{D}$ with partition ratio $\eta=\frac{|\mathcal{T}|}{|\mathcal{S}|}<1$  [^2] The partition ratio $\eta<1$ means that the training set scale is larger than the test set, as a research convention., and $\mathcal{S}_{\mathcal{A}}$, $\mathcal{S}_{\mathcal{P}}$, $\mathcal{T}_{\mathcal{A}}$, $\mathcal{T}_{\mathcal{P}}$, denote the candidate algorithms and problems in $\mathcal{S}$ and $\mathcal{T}$. $L$ is the Lipschitz constants of the model excluding algorithm embedding layer, $\gamma$ is the Lipschitz constant of the loss function, and $W^{(0)}$ is the parameters composed of the problems features and parameters in algorithm embedding layer. $Error_{\mathcal{T}}(\mathbf{f})$ and $Error_{\mathcal{S}}(\mathbf{f})$ denote the test error and empirical error of Modela, and $c_{0}\triangleq\sqrt{\frac{32\ln(4e)}{3}}$. Then with probability of at least $1-\delta$, for all $\ell\left(\mathbf{f}(x_{i}),y_{i}\right)\in\mathcal{L}$,
@@
According to Theorem 2, we observe that the generalization error under the usage of adaptive features is related to three factors: the training data $|\mathcal{S}_{\mathcal{P}}|\cdot|\mathcal{S}_{\mathcal{A}}|$, the partition ratio $\eta$, and some parameters in model including $L$, $W^{(0)}$, and $\gamma$. When the size of the training set is large enough such that $|\mathcal{S}|\gg\frac{\gamma^{2}L^{2}\|W^{(0)}\|^{2}_{2}(1+\eta)^{2}}{c^{2}\eta}$ where $c=c_{0}(1+\eta)+\sqrt{\frac{1}{2}(1+\eta)\ln\frac{1}{\delta}}$, the trend of the generalization error tends to decrease continuously, mainly following the slack term $\mathcal{O}\left(\frac{c}{\sqrt{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$ as the training scale grows. While if the training data is insufficient, the model parameters and the size of the training set will jointly affect the generalization error, mainly following the slack term $\mathcal{O}\left(\frac{\gamma L\|W^{(0)}\|_{2}(1+\eta)}{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}\right)$. This is because the model may fail to capture all the patterns in the data, leading to underfitting. In such cases, it is not advisable to have too many or overly complex model parameters. Based on Theorem 2, it can also be observed that the generalization error is relevant to the product of the norms of each layer’s parameter matrix, as determined by the computation of the upper bound for the Lipschitz constant $L$. Because the $L_{2}$-norm emphasizes the maximum singular value of the parameter matrix [36], outliers in parameter matrix may have a significant impact on the maximum singular value. This highlights the importance of employing preprocessing techniques such as normalization and outlier detection in practical usage to mitigate the occurrence of outliers in the parameter matrices of neural networks. In our subsequent research within the inductive setting, we will derive a bound represented by Frobenius norm, which will demonstrate enhanced resilience to outliers [37], providing better control over the upper bound of generalization error.
@@
On the other hand, it is not difficult to observe from Theorem 2 that, when using the adaptive algorithm feature, the increase in the number of candidate algorithms will have a positive impact on reducing generalization error, following the slack term $\mathcal{O}\left(\sqrt{\frac{c}{\eta|\mathcal{S}_{\mathcal{A}}|\cdot|\mathcal{S}_{\mathcal{P}}|}}\right)$, which is significantly different from traditional learning approaches that focus on learning the mapping from problems to algorithms. When the algorithm feature is absence, for example, some studies model algorithm selection as a multi-class classification problem or performance regression problem based on problem features. In this case, the derivation process is similar to Theorem 2, which allows us to easily obtain upper bounds for the transductive generalization error of the model under the performance regression approach (Modelreg) and the multi-class classification approach (Modelcla). We present them in the form of Corollary 2: